OpenStudy (anonymous):

TRUE OR FALSE. A city has a population of 219,000 citizens. The population decreases by 4% over the course of a year. The new population is 215,040 citizens.

OpenStudy (ash2326):

China gurl do you know how to calculate percentage change?

OpenStudy (anonymous):

no

OpenStudy (ash2326):

okay City has population 219000, let x be the decrease in people We are given 4 % decrease so \[ 4 =\frac{x}{219000} \times 100\] so \[x= \frac{219000\times 4}{100}\] we get x=8760 this is the decrease in population New population= 219000-8760= 210240 so false

OpenStudy (anonymous):

or multiply 219,000 x .96 and get the same answer

OpenStudy (anonymous):

Why don't you learn, here the formula: decrease% = different/ initial = (215,040 - 219,000 ) / 215,040 = - 18.4 %

OpenStudy (anonymous):

my formula is 1.00 = 100% subtract whatever decrease from 1.00 in this case .04 and you get .96 then multiply

OpenStudy (anonymous):

So decrease is NOT 4%

OpenStudy (anonymous):

Oh yeah those numbers don't add up --- where did those 5000 people come from in the original question

OpenStudy (anonymous):

\[219000*(100\%-4\%)\to 219000*0.96=210240 \]